Western Electric has 25,000 shares of common stock outstanding at a price per share of $57 and a rate of return of 14.2 percent. The firm has 7,000 shares of 7 percent preferred stock outstanding at a price of $48 a share. The preferred stock has a par value of $100. The outstanding debt has a total face value of $350,000 and currently sells for 102 percent of face. The yield-to-maturity on the debt is 8.49 percent. What is the firm's weighted average cost of capital if the tax rate is 30 percent?

*NEED IN EXCEL FORMAT*
Answer #1

Please see the table below. Please be guided by the column M to understand the mathematics. The cell highlighted in yellow is your answer. Figures in parenthesis, if any, mean negative values. All financials are in $. Adjacent cells in blue contain the formula in excel I have used to get the final output.

M N S Component Market Value Vi к 13 Common Stock 14 Number 15 Price 16 Market value 17 Cost of equity 25,000 57 1,425,000 =
Answer #2

SOLUTION  (In word document) :


CORREC ANSWER :


EQUITY :


Market value of equity = E = No. of shares * MV per share = 25000 * 57 = 1425000 ($)


Return on equity = rE = 14.20 %


PREFERRED STOCK :


Market value of Preferred stock = PS = No. Of shares * MV = 7000 * 48 = 336000 ($)


Return on PS = rP = 7 / 48 = 0.1458 = 14.58 % 


DEBT :


Market value of debt = D = FV * price per $ of FV = 350000 * 102/100 = 357000 ($) 


Effective interest on debt = rD =  YTM * (1 - T) = 8.49 * (1 - 0.3) = 5.94 % 


TOTAL CAPITA MV BASED :


Total capital as per MV = C = E + PS + D = 1425000 + 336000 + 357000 = 2118000 ($)


WEIGHTS OF CAPITAL COMPONENTS :


Weight of equity = wE = E / C = 1425000 / 2118000 = 0.6728


Weight of PS = wPS = PS / C = 336000 / 2118000 = 0.1586 


Weight of debt = wD = D / C = 357000 / 2118000 = 0.1686 



WEIGHED AVERAGE CAPITAL COST ;


WACC 

= wE * rE + wPS * rPS + wD * rD 

= 0.6728 * 14.20 + 0.1586 * 14.58 + 0.1686 * 5.94

= 12.87 % approx.:   (ANSWER)

answered by: Tulsiram Garg